Grade 12Most Scoring Topics in IIT JEE

The moon revolves round the earth 13 times in one year. if the ratio of the sun-earth distance to the earth-moon distance is 392, then the atio of the masses of sn and earthwill be............

srinivas
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
T²(es) = 4 (pi)² R³(es)/G M(s)
T²(me) = 4 (pi)² R³(me)/G M(e)
[T(me)/T(es)]² = R³(me) M(s)/ R³(es) M(e)
(1/13)² = (1/392)³ M(s)/M(e)
M(s)/M(e) = 392³/13²
Regards
Arun (askIITians forum expert)